carson:

Sheryl purchased a plot of land for $32,500. The land appreciates about 4.9% each year. What is the value of the land after eleven years?

Vocaloid:

final value = A(1+r)^t where A = initial purchase value, rate = 4.9/100, and t = 11 yrs

Vocaloid:

@carson still there?

carson:

i figured it out shortly after i posted this but thank you for responding so quickly @Vocaloid